maths89898:

(Posting a question for iSuckAtmath101 because their questions won't work) A roller coaster’s height is given by the equation h = –.086t^2 + 6t + 50, where t represents the time in seconds. How long will it take riders to pass over the hill and reach ground level? hint: Set h = 0. 34.88 seconds 50.00 seconds 50.17 seconds 77.29 seconds

maths89898:

@isuckatmath101

maths89898:

@jhonyy9 can you help math with this?

surjithayer:

it will touch the ground when h=0 \[-086t^2+6t+50=0\] or \[0.043t^2-3t-25=0\] \[t=\frac{ 3\pm \sqrt{9-4\times 0.043\times(-25)} }{ 2\times 0.043 }\] \[t=\frac{ 3\pm \sqrt{9+4.3} }{ 0.086 }\] \[t=\frac{ 3\pm \sqrt{13.3} }{ 0.086 }\] \[t=\frac{ 3\pm 3. 646917 }{ 0.086 }=\frac{ -0.646917 }{ 0.086 },\frac{ 6.646917 }{ 0.086 }\] t=--7.52,77.29 rejecting negative sign time of flight=77.29 seconds.

maths89898:

@surjithayer wrote:
it will touch the ground when h=0 \[-086t^2+6t+50=0\] or \[0.043t^2-3t-25=0\] \[t=\frac{ 3\pm \sqrt{9-4\times 0.043\times(-25)} }{ 2\times 0.043 }\] \[t=\frac{ 3\pm \sqrt{9+4.3} }{ 0.086 }\] \[t=\frac{ 3\pm \sqrt{13.3} }{ 0.086 }\] \[t=\frac{ 3\pm 3. 646917 }{ 0.086 }=\frac{ -0.646917 }{ 0.086 },\frac{ 6.646917 }{ 0.086 }\] t=--7.52,77.29 rejecting negative sign time of flight=77.29 seconds.
man sorry for closing the question on you but good job